OpenStudy (calculusxy):

MEDAL FOR ANSWER! Esteban is making turtle brownies. The recipe calls for 3/4 bag of caramel squares. One bag has 24 caramel squares in it. a. Esteban decides to make two batches of turtle brownies. Write a number sentence to show how many bags of caramel squares he will use. @mathstudent55 @phi

OpenStudy (anonymous):

Esteban will use double the amount shown in the recipe which is six over four and this means one and a half bags of caramel squares which is Thirty Six caramel squares

OpenStudy (calculusxy):

Is this correct. @phi ? Because I got the same answer when doing it, but I am really not sure.

OpenStudy (phi):

what answer did you get ?

OpenStudy (calculusxy):

I got thirty six.

OpenStudy (calculusxy):

@phi

OpenStudy (phi):

OK, but the question is *** Write a number sentence to show how many bags of caramel squares he will use. *** how many bags ?

OpenStudy (calculusxy):

So do I divide 36 by 3/4?

OpenStudy (phi):

No. think like this: two batches will take 2 times more than one batch one batch needs ¾ of a bag two batches will need how many bags?

OpenStudy (calculusxy):

6/8?

OpenStudy (phi):

2*¾ is not 6/8 when you multiply fractions, you multiply top times top and bottom times bottom. here, do \[ \frac{2}{1} \cdot \frac{3}{4} \]

OpenStudy (calculusxy):

Yes. I got mixed up with finding equiv. fractions. But the 2 x 3/4 is 6/4.

OpenStudy (phi):

which you can simplify, right ?

OpenStudy (calculusxy):

Yes. As a mixed number it would be. 1 1/2 and as a fraction it can be 3/2.

OpenStudy (calculusxy):

@phi What would be my answer to the question. I don't mean you saying the answer to me directly. Help will be fine.

OpenStudy (phi):

Write a number sentence to show how many bags of caramel squares he will use. I think they want you to write down 2 * ¾ = 1 ½

OpenStudy (calculusxy):

So it's 1 1/2 batches then?

OpenStudy (phi):

The answer is not just a number. They want a "number sentence" which is goofy talk for writing down the math you use to find 1 ½

OpenStudy (calculusxy):

Ok thank you. Can you help me with one more question?

OpenStudy (phi):

If you make it a new post.